calculate half-life using volume of distribution of 462L and clearance of 40mL/min
Answer #1

ANSWER.

* Calculation of half-life.

       Formula

             Half life (t1/2) = 0.693 * Volume of distribution(Vd) / Clearance(CL).

Here

        Volume of distribution(Vd) = 462L.

Convert litre to millilitre for calculation

             1 l = 1000 ml

So

       Volume of distribution (Vd) = 462 * 1000

                                                  = 462000 ml.

             Clearance (CL) = 40 mL/min.

As per above formula

          Half life(t1/2) = 0.693 * 462000/40

                               = 0.693 * 11550

                               = 8004 min.

             Convert min to hr by divide with 60

      So

              Half life (t1/2) = 8004/60

                                    = 133.4 hr